# Sign problem in tensor network contraction

### v1

@@ Acknowledgements.
We acknowledge helpful conversations with Garnet Chan, Ryan Levy, Daniel Malz, David Perez-Garcia, Bram Vanhecke, and Frank Verstraete. This research was funded in part by the Austrian Science Fund FWF (Grant DOIs 10.55776/COE1, 10.55776/P36305, and 10.55776/F71), the European Union – NextGenerationEU, and the European Union’s Horizon 2020 research and innovation programme through Grant No. 863476 (ERC-CoG SEQUAM). Jiaqing Jiang is supported by MURI Grant FA9550-18-1-0161 and the IQIM, an NSF Physics Frontiers Center (NSF Grant PHY-1125565). D.H. acknowledges financial support from the US DoD through a QuICS Hartree fellowship. Part of this work was conducted while the authors were visiting the Simons Institute for the Theory of Computing during summer 2023 and spring 2024, whose hospitality we gratefully acknowledge. The computational results have been in part achieved using the Vienna Scientific Cluster (VSC). For open access purposes, the authors have applied a CC BY public copyright license to any author accepted manuscript version arising from this submission.

### v2

@@ Acknowledgements.
We acknowledge helpful conversations with Garnet Chan, Ryan Levy, Daniel Malz, David Perez-Garcia, Bram Vanhecke, and Frank Verstraete. We would also like to thank Paul Brehmer and Anna Francuz for providing us with the PEPS tensor for the $J_{1}$-$J_{2}$ model. This research was funded in part by the Austrian Science Fund FWF (Grant DOIs 10.55776/COE1, 10.55776/P36305, and 10.55776/F71), the European Union – NextGenerationEU, and the European Union’s Horizon 2020 research and innovation programme through Grant No. 863476 (ERC-CoG SEQUAM). Jiaqing Jiang is supported by MURI Grant FA9550-18-1-0161 and the IQIM, an NSF Physics Frontiers Center (NSF Grant PHY-1125565). D.H. acknowledges financial support from the US DoD through a QuICS Hartree fellowship. Jielun Chen is supported by the US National Science Foundation under grant No. CHE-2102505. Part of this work was conducted while the authors were visiting the Simons Institute for the Theory of Computing (supported by DOE QSA grant #FP00010905) during summer 2023 and spring 2024, whose hospitality we gratefully acknowledge. The computational results have been in part achieved using the Vienna Scientific Cluster (VSC). For open access purposes, the authors have applied a CC BY public copyright license to any author accepted manuscript version arising from this submission.